IF THE DISTANCE BETWEEN THE EARTH AND THE SUN WERE HALF ITS PRESENT VALUE THE NUMBER OF DAYS IN A YEAR WOULD HAVE BEEN

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
 

The ratio of the time periods would be {new is say T1, actual is say T2}

 

 

 

Note that T is directly proportional to √(R3)

 

 

 

T1/T2 = √([R/2]3)/√(R3) = 1/(2√2)

 

 

 

Hence T1 = T2/2√2 = 365/2√2

 

 

 

Which is approx 129 days.

 

 

Regards

Arun (askIITians forum expert)